OpenStudy (anonymous):

Debbie and Trisha are looking at the equation the square root of the quantity of 4 times x minus 6 equals square root of x . Debbie says that the solution is extraneous. Trisha says that the solution is non-extraneous. Is Debbie correct? Is Trisha correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (anonymous):

@@Zarkon

OpenStudy (anonymous):

@Zarkon

OpenStudy (mathmale):

Much depends on whether we're interpreting the verbal description of the equation correctly. My interpretation is\[\sqrt{4x-6}=\sqrt{x}\] Does that agree with yours?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Then please square both sides of this equation, separately. Type the result here.

OpenStudy (anonymous):

i really dont know @mathmale i need a walk through

OpenStudy (anonymous):

@micahpkay

OpenStudy (anonymous):

@kawaiicat123

OpenStudy (anonymous):

@charlotte123

OpenStudy (anonymous):

@charlotte123